OpenStudy (benzo):

h5p2.....im lost on this....give me step by step how you do this please....What is the minimum value of VDD (in Volts) that we must supply to keep the transistor in the region of saturated operation?

OpenStudy (anonymous):

what are ur values given.............give the cumplete question

OpenStudy (anonymous):

dude answer something between 2 to 7

OpenStudy (anonymous):

try it

OpenStudy (anonymous):

i guess you got it

OpenStudy (benzo):

not yet

OpenStudy (anonymous):

vdd=5.5. just try it

OpenStudy (benzo):

i got it...VIN-VOUT greater or equall ti VT...VT=2.0 6.1-3.6=2.5 5.4-2.9=2.5 6.8-4.2=2.6 VIN-VT 6.8-2=4.8 VDD=4.8

OpenStudy (anonymous):

is you check. i though it must be corect

OpenStudy (anonymous):

vdd is 5.5 for me. other vin is diferent

OpenStudy (benzo):

ok right for me....thanks.....H5P1.....whats the step by step solution 1, 2, 3

OpenStudy (anonymous):

-.5,8000,0.093 is the answere

OpenStudy (benzo):

step by step solution

OpenStudy (anonymous):

, we have to find max Vin ; to find Vin let us apply kvll to the input loop Vin - Vgs+1 =0; ie, Vin = Vgs-1 ...............................(eqn1) now the unknown parameter is Vgs, hence to find it lets apply KVL to external loop : 1-Ids*R-Vds+1=0 , further substituting the value of Ids = k/2 *(Vgs-Vt)^2 in the above equation and solving the quadratic equation we will get, Vds = 0.593, -0.84 : as Vds needs tobe +ve hence neglecting he -ve term we have Vds = 0.593 and we know Vds = Vgs-Vt ie Vgs = Vds+Vt therefore, Vgs = 0.593+0.5=1.093 and substituting the above valuw in eqn1 we have Vin=1.093-1=0.093

OpenStudy (anonymous):

@benzo : In saturation: iDS=K/2*(vGS-VT)2, and since here vGS is vIN-vOUT, then iDS=K/2*(vIN-vOUT-VT)2 Q2 vOUT=

OpenStudy (anonymous):

vOUT=iDS*RS

OpenStudy (benzo):

thanks lab and hw 100%

OpenStudy (anonymous):

i get the vout for sorce flower small signal but its doesn't work what you vout for source flower small signal

OpenStudy (anonymous):

vi=0.0009V what is the incremental output vo? also Gain

OpenStudy (anonymous):

gain=0.9498

OpenStudy (anonymous):

and vo?

OpenStudy (anonymous):

yes vo=0.9498

OpenStudy (anonymous):

Wont that conclude a gain of 1055?

OpenStudy (anonymous):

Bro Gain is .94 vo=?

OpenStudy (anonymous):

Q1 vOUT = iDS*RS vout/vin = ϑvOUT/ϑvIN @ VIN = RS * ϑiDS/ϑvIN = RS * (1/RS + 1/(K*RS 2)*(-2K*RS/(2d)) where d=sqrt(1+2K*RS*(VIN-VT))) Simplifying the expression above will give: vout/vin = (1-1/d) à vout=vin*(1-1/sqrt(1+2K*RS*(VIN-VT)))) Q2 For K=2 [A/V2], VT=2V, RS=16Ω, VIN=6.5V and vin=0.001 V, d=19.925 and thus vout/vin = (1-1/d) à vout=0.001*(1-1/19.925)=0.001*0.9498 à vout = 0.0009498 V Q3 We've already found voltage gain solving for Q2. vout/vin = (1-1/d) = 1-1/19.925 = 0.9498 You can see that source follower has a voltage gain slightly less than 1.

OpenStudy (anonymous):

vo= 0.0009498 gives X (wrong)

OpenStudy (anonymous):

for each value each answer

OpenStudy (anonymous):

for VIN = 6.4, 6.9 and 5.9?